OpenStudy (anonymous):

Carson found an old tent in the attic of his house and decided to set it up in the back yard. However, the support sticks for the tent are missing. If the tent is 72 inches across on the bottom and 45 inches on each side, how tall of a stick does he need to set up the tent?

OpenStudy (anonymous):

(72/2)^2 + b^2 = 45^2 36^2 + b^2 = 45^2 1296 + b^2 = 2025 b^2 = 729 b = 27 inches